The directions were to...
Use a system of linear equations to solve each problem.
1. Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test.How many of each type of question are on the test?

Answer by stanbon(75887)

Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point and four-point questions on the test.How many of each type of question are on the test?
---------
Let number of 2-point tests be "x"; These are worth 2x points
Number of 4-point tests is "40-x"; These are worth 4(40-x)=160-4x points
-----------
EQUATION:
points + points = 100
2x + 160-4x = 100
-2x = -60
x=30 (number of 2-pt tests)
40-x=10 (number of 4-pt tests)
Cheers,
Stan H.